OpenStudy (gabylovesyou):

@hartnn John and Tim are looking at the equation the square root of the quantity of 3 times x minus 4 equals square root of x . John says that the solution is extraneous. Tim says the solution is non-extraneous. Is John correct? Is Tim correct? Are they both correct? Justify your response by solving this equation, explaining each step with complete sentences.

OpenStudy (gabylovesyou):

\[\sqrt{3x - 4} = \sqrt{x}\]

OpenStudy (gabylovesyou):

@hartnn

OpenStudy (gabylovesyou):

@sleepyjess

OpenStudy (gabylovesyou):

@ash2326

OpenStudy (sleepyjess):

We need to solve for x. So simply get rid of the square roots first.

OpenStudy (gabylovesyou):

so 3x/3 - 4 = x - 3 4 = x - 3 ?

OpenStudy (sleepyjess):

How are you getting that?

OpenStudy (gabylovesyou):

oh nvm cx lmao

OpenStudy (sleepyjess):

okay haha

OpenStudy (sleepyjess):

do you know how to get rid of the square root signs?